A PIR motion detector made with an infrared faceted fresnel lens has the following properties:

  • Focal length of single facet lens = 22.5 mm
  • Distance of object from lens = 12.5 m
  • Minimum displacement of object which results in detection = 250 mm
  • Number of facets = 8

What is the facet pitch in mm?


1
Expert's answer
2021-04-01T01:04:25-0400

Focal length of single facet lens = 22.5 mm

Distance of object from lens = 12.5 m

Minimum displacement of object which results in detection = 250 mm

Number of facets = 8


distance from minimum displacement = 12.5m - 250mm = 12.25m


"\dfrac{1}{T}" = "\dfrac1{12250} - \dfrac1{22.5} = \dfrac{4890}{110250}"


T = 110250/4890 = 22.54mm


facet pitch = 22.54mm/8 = 2.82mm